# Testing Conditional Independence via the Spectral Generalized Covariance Measure: Beyond Euclidean Data

### v2

@@ Appendix A Proofs of Main Theorems of the conditional cross-covariance operator
By (A.2), we have $\Sigma_{\ddot{X}Y\mid Z}=0$. Under Assumption 2.2 (i), both $k_{\ddot{\mathcal{X}}}=k_{\mathcal{X}}\otimes k_{\mathcal{Z}}$ and $k_{\mathcal{Y}}$ are characteristic. Fix arbitrary $f\in L^{2}_{XZ}$ and $g\in L^{2}_{Y}$. Because each kernel is bounded and characteristic by Assumptions 2.1 and 2.2 (i), the direct sum $\mathcal{H}_{\ddot{\mathcal{X}}}+\mathbb{R}$ is dense in $L^{2}(P_{XZ})$ for any probability measure $P_{XZ}$ on $(\ddot{\mathcal{X}},\mathcal{B}(\ddot{\mathcal{X}}))$, and similarly $\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$ is dense in $L^{2}(P_{Y})$ for any probability measure $P_{Y}$ on $(\mathcal{Y},\mathcal{B}(\mathcal{Y}))$ (Fukumizu, Bach and Jordan, 2009, Proposition 5).
@@
Using the Cauchy-Schwarz inequality and the fact that conditional expectation is an $L^{2}$-contraction, we obtain
$\left|B(f,g)\right|\leq\|f\|_{L_{XZ}^{2}}\|g\|_{L_{Y}^{2}},$
@@
Next, for any $f_{H}\in\mathcal{H}_{\ddot{\mathcal{X}}}$ and $g_{H}\in\mathcal{H}_{\ddot{\mathcal{Y}}}$, expanding the definition of $\Sigma_{\ddot{X}Y\mid Z}$ yields
@@
$B(f_{H},g_{H})=0\quad\forall f_{H}\in\mathcal{H}_{\ddot{\mathcal{X}}},\ g_{H}\in\mathcal{H}_{\ddot{\mathcal{Y}}}.$
Moreover, constants vanish under centering, so $B$ is identically zero on $(\mathcal{H}_{\ddot{\mathcal{X}}}+\mathbb{R})\times(\mathcal{H}_{\ddot{\mathcal{Y}}}+\mathbb{R})$.
@@
$M_{f,g}(Z)\coloneqq\operatorname{\mathbb{E}}\left[\tilde{f}\,\tilde{g}\mid Z\right]\in L^{2}\left(P_{Z}\right).$
@@
$\bigl\langle M_{f_{H},g_{H}},\,h_{H}\bigr\rangle_{L^{2}(P_{Z})}=0\quad\text{for all }h_{H}\in\mathcal{H}_{\mathcal{Z}}.$
Since $\mathcal{H}_{\mathcal{Z}}$ is dense in $L^{2}(P_{Z})$ by Since $\mathcal{H}_{\mathcal{Z}}$ is dense in $L^{2}(P_{Z})$ by Assumption 2.2 (iii), we conclude that
$M_{f_{H},g_{H}}=0\quad\text{in }L^{2}(P_{Z})\quad\text{for all }f_{H}\in\mathcal{H}_{\mathcal{X}}+\mathbb{R},g_{H}\in\mathcal{H}_{\mathcal{Y}}+\mathbb{R}.$
by the self-duality via the $L^{2}$ inner product.
@@
Hence there exist sequences $f_{n}\in\mathcal{H}_{\mathcal{X}}+\mathbb{R}$ and $g_{n}\in\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$ with $\|f_{n}-f\|_{L_{X}^{2}}\to 0$ and $\|g_{n}-g\|_{L_{Y}^{2}}\to 0$. Let $\tilde{f}_{n}=f_{n}(X)-\operatorname{\mathbb{E}}[f_{n}(X)\mid Z]$ and $\tilde{g}_{n}=g_{n}(Y)-\operatorname{\mathbb{E}}[g_{n}(Y)\mid Z]$. Since conditional expectation is an $L^{2}$-contraction, $\|\tilde{f}_{n}-\tilde{f}\|_{L_{X}^{2}}\leq 2\|f_{n}-f\|_{L_{X}^{2}}\to 0$ and $\|\tilde{g}_{n}-\tilde{g}\|_{L_{Y}^{2}}\to 0$. Therefore
@@ B.2 Uniformly Asymptotic Properties
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\hat{C}_{n}-C_{P}\right\|_{\operatorname{tr}}>c_{0}/2\right)\xrightarrow[n\to\infty]{}0.$ (A.13)
@@
Proof of Lemma B.2.
First, on $\mathcal{E}_{n}(P)$ one has
$Z_{n}\geq\lambda_{\max}(\hat{C}_{n})\,\chi^{2}_{1}\geq\frac{\lambda_{\star}}{2}\,\chi^{2}_{1},\qquad\operatorname{\mathbb{E}}\left[Z_{n}\right]=\operatorname{tr}(\hat{C}_{n})\leq 2c_{1}.$
Hence, by monotonicity of quantiles and Markov’s inequality,
$Q_{\tau}(\hat{C}_{n})\geq\frac{\lambda_{\star}}{2}\,Q_{\tau}\left(\chi^{2}_{1}\right)=a_{\mathcal{T}},\qquad Q_{\tau}(\hat{C}_{n})\leq\frac{2c_{1}}{1-\tau}\leq\frac{2c_{1}}{1-\tau_{+}}=b_{\mathcal{T}}.$
Write $Z_{n}=U_{s}+R$ with $S=\{1,\dots,K(\eta_{\mathcal{T}})\}$, where $U_{s}\coloneqq\sum_{k\in S}\hat{\lambda}_{k}\,\chi^{2}_{1,k}$ and $R\coloneqq\sum_{k\notin S}\hat{\lambda}_{k}\,\chi^{2}_{1,k}$ are independent. On $\mathcal{E}_{n}(P)$,
@@
$\operatorname{\mathbb{E}}\left[R\right]=\sum_{k\notin S}\hat{\lambda}_{k}\leq\frac{a_{\mathcal{T}}}{8}.$
For $r_{0}\coloneqq a_{\mathcal{T}}/4$, Markov’s inequality yields
$\mathbb{P}\left(R\leq r_{0}\right)\geq 1-\frac{\operatorname{\mathbb{E}}[R]}{r_{0}}\geq\frac{1}{2}.$
We first note the following positivity, which will later yield a uniform lower bound on the compact strip once continuity is established. For any $x>0$ and $\lambda\in\Lambda_{S}$ with $\hat{\lambda}_{1}\geq\lambda_{\star}/2>0$, the convolution representation of the law of $U_{S}=\sum_{k\in S}\hat{\lambda}_{k}\chi^{2}_{1,k}$ implies $g_{S}(x;\lambda)>0$ because the kernel $h_{\hat{\lambda}_{1}}(\cdot)$ is strictly positive on $(0,\infty)$ and the remaining coordinates contribute a nonnegative mixing measure (allowing atoms at 0 when some $\hat{\lambda}_{k}=0$). Hence, once continuity on the compact set $I_{x}\times\Lambda_{S}$ is proved below, the extreme value theorem gives
$m_{S}\coloneqq\inf_{\lambda\in\Lambda_{S}}\ \inf_{x\in I_{x}}g_{S}(x;\lambda)\ >\ 0.$
For $\hat{\lambda}>0$ set
$h_{\hat{\lambda}}(u)\coloneqq\frac{1}{\hat{\lambda}}\,f_{\chi^{2}_{1}}\left(\frac{u}{\hat{\lambda}}\right)\coloneqq\frac{1}{\sqrt{2\pi\hat{\lambda}u}}\exp\left(-\frac{u}{2\hat{\lambda}}\right),\qquad u>0,$ (A.17)
and define $h_{0}(u)\equiv 0$ for $u>0$. Let $I_{x}\coloneqq\left[\frac{a_{\mathcal{T}}}{4},\,2b_{\mathcal{T}}\right]\subset(0,\infty)$ and
$\Lambda_{s}\coloneqq\left\{\lambda=(\hat{\lambda}_{k})_{k\in S}\colon\hat{\lambda}_{1}\in\left[\frac{\lambda_{\star}}{2},\,2c_{1}\right],\ \ 0\leq\hat{\lambda}_{k}\leq 2c_{1},\ \ \sum_{k\in S}\hat{\lambda}_{k}\leq 2c_{1}\right\}.$
For $x>0$ and $\lambda\in\Lambda_{S}$, the density of $U_{S}$ can be written as
$g_{S}(x;\lambda)=\int_{(0,\infty)^{|S|-1}}h_{\hat{\lambda}_{1}}\left(x-\sum_{j=2}^{|S|}t_{j}\right)\,\prod_{j=2}^{|S|}h_{\hat{\lambda}_{j}}(t_{j})\,\mathds{1}_{\left\{\sum_{j=2}^{|S|}t_{j}<x\right\}}\mathop{}\!\mathrm{d}t_{2}\cdots\mathop{}\!\mathrm{d}t_{|S|}.$ (A.18)
Let $(x_{n},\lambda^{(n)})\to(x,\lambda)$ in $I_{x}\times\Lambda_{S}$ and define, for $t=(t_{2},\dots,t_{|S|})$,
$\Phi_{x,\lambda}(t)\coloneqq h_{\hat{\lambda}_{1}}\left(x-\sum_{j=2}^{|S|}t_{j}\right)\,\prod_{j=2}^{|S|}h_{\hat{\lambda}_{j}}(t_{j})\,\mathds{1}_{\left\{\sum_{j=2}^{|S|}t_{j}<x\right\}}(t).$
If $\sum_{j=2}^{|S|}t_{j}\neq x$, then $(x,\lambda)\mapsto\Phi_{x,\lambda}(t)$ is continuous: for fixed $t_{j}>0$, (A.17) shows continuity of $\hat{\lambda}_{j}\mapsto h_{\hat{\lambda}_{j}}(t_{j})$ on $0,2c_{1}]$ with $h_{\hat{\lambda}_{j}}(t_{j})\to 0=h_{0}(t_{j})$ as $\hat{\lambda}_{j}\downarrow 0$, and $u\mapsto h_{\hat{\lambda}_{1}}(u)$ is continuous on $(0,\infty)$. Since the boundary $\{\sum_{j=2}^{|S|}t_{j}=x\}$ is Lebesgue-null, we have
$\Phi_{x_{n},\lambda^{(n)}}(t)\ \longrightarrow\ \Phi_{x,\lambda}(t)\qquad\text{for a.e.\ }t.$
To obtain a parameter-free integrable envelope, note from (A.17) that there exists $C=C(\lambda_{\star},c_{1})$ such that
$h_{\hat{\lambda}_{1}}(u)\leq Cu^{-1/2}\kern 5.0pt(u>0,\hat{\lambda}_{1}\geq\lambda_{\star}/2),\quad h_{\hat{\lambda}_{j}}(t_{j})\leq Ct_{j}^{-1/2}\kern 5.0pt(t_{j}>0,2\leq j\leq|S|,\hat{\lambda}_{j}\leq 2c_{1}).$
Since $x\leq 2b_{\mathcal{T}}$ for $x\in I_{x}$, it follows that
$\Phi_{x,\lambda}(t)\leq C\,\left(2b_{\mathcal{T}}-\sum_{j=2}^{|S|}t_{j}\right)_{+}^{-1/2}\prod_{j=2}^{|S|}t_{j}^{-1/2}\,\mathds{1}_{\left\{\sum_{j=2}^{|S|}t_{j}<2b_{\mathcal{T}}\right\}}(t)\eqqcolon C\,M(t).$
Here, $\int M(t)\mathop{}\!\mathrm{d}t<\infty$ holds because of Dirichlet–Beta formula.
By the dominated convergence theorem applied to (A.18) with the a.e. convergence above and the envelope CMC\,M, we obtain
$g_{S}(x_{n};\lambda^{(n)})\ \longrightarrow\ g_{S}(x;\lambda)\qquad\text{as }(x_{n},\lambda^{(n)})\to(x,\lambda)\ \text{ in }I_{x}\times\Lambda_{S}.$
Therefore $(x,\lambda)\mapsto g_{S}(x;\lambda)$ is continuous on $I_{x}\times\Lambda_{S}$. Combining this continuity with the initial positivity and the compactness of $I_{x}\times\Lambda_{S}$ yields the uniform density lower bound $m_{S}>0$ asserted at the beginning.
Moreover, since $\hat{\lambda}_{1}\geq\lambda_{\star}/2>0$, each $h_{\hat{\lambda}_{1}}$ is strictly positive on $(0,\infty)$; convolution with nonnegative measures preserves strict positivity on $(0,\infty)$, hence $g_{S}(x;\lambda)>0$ for all $x>0$ and all $\lambda\in\Lambda_{S}$. Therefore, by the extreme value theorem on the compact set $a_{\mathcal{T}}/4,\,2b_{\mathcal{T}}]\times\Lambda_{S}$,
$m_{s}\coloneqq\inf_{\lambda\in\Lambda_{s}}\ \inf_{x\in[a_{\mathcal{T}}/4,\,2b_{\mathcal{T}}]}\ g_{s}(x;\lambda)\ >\ 0.$
Let $f_{n}$ be the density of $Z_{n}$. For any $x\in\left[a_{\mathcal{T}}/2,2b_{\mathcal{T}}\right]$, since $x-v\in\left[a_{\mathcal{T}}/4,2b_{\mathcal{T}}\right]$ whenever $v\in[0,r_{0}]$, one has
$f_{n}(x)$ $=\int_{[0,\infty)}g_{s}(x-v)\mathop{}\!\mathrm{d}F_{R}(v)$
$\geq\int_{[0,r_{0}]}g_{s}(x-v)\mathop{}\!\mathrm{d}F_{R}(v)$
$\geq m_{s}\,\mathbb{P}_{P}\left(R\leq r_{0}\right)$
$\geq\frac{m_{s}}{2}\eqqcolon m_{0}>0.$
Since $Q_{\tau}(\hat{C}_{n})\in\left[a_{\mathcal{T}},\,b_{\mathcal{T}}\right]$, one has $Q_{\tau}(\hat{C}_{n})-r_{0}\geq a_{\mathcal{T}}-a_{\mathcal{T}}/4=a_{\mathcal{T}}/2$ and $Q_{\tau}(\hat{C}_{n})+r_{0}\leq b_{\mathcal{T}}+a_{\mathcal{T}}/4\leq 2b_{\mathcal{T}}$. Therefore $Q_{\tau}(\hat{C}_{n})-r_{0},\,Q_{\tau}(\hat{C}_{n})+r_{0}\,]\subset[a_{\mathcal{T}}/2,\,2b_{\mathcal{T}}]$, and the displayed lower bound for $f_{n}$ holds on this interval, as claimed. ∎

### v3

@@ Appendix A Proofs of Main Theorems of the conditional cross-covariance operator
By (A.2), we have $\Sigma_{\ddot{X}Y\mid Z}=0$. Under Assumption 2.2 (i), both $k_{\ddot{\mathcal{X}}}=k_{\mathcal{X}}\otimes k_{\mathcal{Z}}$ and $k_{\mathcal{Y}}$ are characteristic. Because each kernel is bounded and characteristic by Assumptions 2.1 and 2.2 (i), the direct sum $\mathcal{H}_{\ddot{\mathcal{X}}}+\mathbb{R}$ is dense in $L^{2}(P_{XZ})$ for any probability measure $P_{XZ}$ on $(\ddot{\mathcal{X}},\mathcal{B}(\ddot{\mathcal{X}}))$, and similarly $\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$ is dense in $L^{2}(P_{Y})$ for any probability measure $P_{Y}$ on $(\mathcal{Y},\mathcal{B}(\mathcal{Y}))$ (Fukumizu, Bach and Jordan, 2009, Proposition 5).
@@
Using the Cauchy–Schwarz inequality and that conditional expectation is an $L^{2}$-contraction, we obtain
$|B(f,g)|\leq\|f-\operatorname{\mathbb{E}}[f\mid Z]\|_{L^{2}}\,\|g-\operatorname{\mathbb{E}}[g\mid Z]\|_{L^{2}}\leq 4\|f\|_{L_{XZ}^{2}}\|g\|_{L_{Y}^{2}},$
@@
Next, for any $f_{H}\in\mathcal{H}_{\ddot{\mathcal{X}}}$ and $g_{H}\in\mathcal{H}_{\mathcal{Y}}$, expanding the definition of $\Sigma_{\ddot{X}Y\mid Z}$ yields
@@
$B(f_{H},g_{H})=0\quad\forall f_{H}\in\mathcal{H}_{\ddot{\mathcal{X}}},\ g_{H}\in\mathcal{H}_{\mathcal{Y}}.$
Moreover, constants vanish under centering, so $B$ is identically zero on $(\mathcal{H}_{\ddot{\mathcal{X}}}+\mathbb{R})\times(\mathcal{H}_{\mathcal{Y}}+\mathbb{R})$.
@@
$M_{f,g}(Z)\coloneqq\operatorname{\mathbb{E}}[\tilde{f}\,\tilde{g}\mid Z]\in L^{1}(P_{Z}).$
@@
For such $f_{H}\in\mathcal{H}_{\mathcal{X}}+\mathbb{R}$ and $g_{H}\in\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$, bounded kernels imply that $f_{H}(X)$ and $g_{H}(Y)$ are essentially bounded; hence $\tilde{f}_{H}$ and $\tilde{g}_{H}$ are essentially bounded as well, so $\tilde{f}_{H}\,\tilde{g}_{H}\in L^{2}(\mathbb{P})$ and therefore
$M_{f_{H},g_{H}}(Z)\coloneqq\operatorname{\mathbb{E}}\left[\tilde{f}_{H}\,\tilde{g}_{H}\mid Z\right]\in L^{2}(P_{Z}).$
@@
$\left\langle M_{f_{H},g_{H}},\,h_{H}\right\rangle_{L^{2}(P_{Z})}=0\quad\text{for all }h_{H}\in\mathcal{H}_{\mathcal{Z}}.$
Since $\mathcal{H}_{\mathcal{Z}}$ is dense in $L^{2}(P_{Z})$ by Assumption 2.2 (iii), the identity
$\langle M_{f_{H},g_{H}},h_{H}\rangle_{L^{2}(P_{Z})}=0\quad\forall\,h_{H}\in\mathcal{H}_{\mathcal{Z}}$
implies $M_{f_{H},g_{H}}=0$ in $L^{2}(P_{Z})$ by the self-duality of $L^{2}(P_{Z})$.
@@
Hence there exist sequences $f_{n}\in\mathcal{H}_{\mathcal{X}}+\mathbb{R}$ and $g_{n}\in\mathcal{H}_{\mathcal{Y}}+\mathbb{R}$ with $\|f_{n}-f\|_{L_{X}^{2}}\to 0$ and $\|g_{n}-g\|_{L_{Y}^{2}}\to 0$. Let $\tilde{f}_{n}=f_{n}(X)-\operatorname{\mathbb{E}}[f_{n}(X)\mid Z]$ and $\tilde{g}_{n}=g_{n}(Y)-\operatorname{\mathbb{E}}[g_{n}(Y)\mid Z]$. Since conditional expectation is an $L^{2}$-contraction,
$\|\tilde{f}_{n}-\tilde{f}\|_{L^{2}}\leq 2\|f_{n}-f\|_{L_{X}^{2}}\to 0,\quad\|\tilde{g}_{n}-\tilde{g}\|_{L^{2}}\leq 2\|g_{n}-g\|_{L_{Y}^{2}}\to 0.$
Therefore
@@ B.2 Uniformly Asymptotic Properties
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\hat{C}_{n}-C_{P}\right\|_{\operatorname{tr}}>\varepsilon\right)\xrightarrow[n\to\infty]{}0\qquad\text{for all }\varepsilon>0.$ (A.13)
@@
Proof.
Fix $P\in\tilde{\mathcal{P}}$ and work on the event $\mathcal{E}_{n}(P)$. Since $\|\cdot\|_{\mathrm{op}}\leq\|\cdot\|_{\mathrm{tr}}$ and $\eta_{\mathcal{T}}\leq\lambda_{\star}/4$, we have
$\|\hat{C}_{n}-C_{P}\|_{\mathrm{op}}\leq\eta_{\mathcal{T}}\leq\frac{\lambda_{\star}}{4}.$
By Weyl’s inequality for self-adjoint operators,
$\lambda_{\max}(\hat{C}_{n})\geq\lambda_{\max}(C_{P})-\|\hat{C}_{n}-C_{P}\|_{\mathrm{op}}\geq\lambda_{\star}-\frac{\lambda_{\star}}{4}\geq\frac{\lambda_{\star}}{2}.$
Moreover, using $|\operatorname{tr}(A)-\operatorname{tr}(B)|\leq\|A-B\|_{\mathrm{tr}}$ and $\eta_{\mathcal{T}}\leq c_{1}/2$,
$\operatorname{tr}(\hat{C}_{n})\leq\operatorname{tr}(C_{P})+\|\hat{C}_{n}-C_{P}\|_{\mathrm{tr}}\leq c_{1}+\frac{c_{1}}{2}\leq 2c_{1}.$
Let $(\hat{\lambda}_{k})_{k\geq 1}$ denote the eigenvalues of $\hat{C}_{n}$ in non-increasing order, and let $(\chi^{2}_{1,k})_{k\geq 1}$ be i.i.d. $\chi^{2}_{1}$ random variables, independent of $\hat{C}_{n}$. Conditionally on $\hat{C}_{n}$, the random variable
$Z_{n}\coloneqq\sum_{k\geq 1}\hat{\lambda}_{k}\,\chi^{2}_{1,k}$
satisfies $Z_{n}\geq\hat{\lambda}_{1}\chi^{2}_{1,1}=\lambda_{\max}(\hat{C}_{n})\chi^{2}_{1,1}$ almost surely. Hence, for any $\tau\in\mathcal{T}$, by monotonicity of quantiles and the scaling property $Q_{\tau}(aX)=aQ_{\tau}(X)$ for $a>0$,
$Q_{\tau}(\hat{C}_{n})=Q_{\tau}(Z_{n}\mid\hat{C}_{n})\geq\frac{\lambda_{\star}}{2}\,Q_{\tau}(\chi^{2}_{1})\geq\frac{\lambda_{\star}}{2}\,Q_{\tau_{-}}(\chi^{2}_{1})=a_{\mathcal{T}}.$
On the other hand, $Z_{n}\geq 0$ and $\operatorname{\mathbb{E}}[Z_{n}\mid\hat{C}_{n}]=\operatorname{tr}(\hat{C}_{n})\leq 2c_{1}$, so Markov’s inequality implies
$Q_{\tau}(\hat{C}_{n})\leq\frac{\operatorname{\mathbb{E}}[Z_{n}\mid\hat{C}_{n}]}{1-\tau}\leq\frac{2c_{1}}{1-\tau}\leq\frac{2c_{1}}{1-\tau_{+}}=b_{\mathcal{T}}.$
Therefore $Q_{\tau}(\hat{C}_{n})\in[a_{\mathcal{T}},b_{\mathcal{T}}]$ for all $\tau\in\mathcal{T}$.
Let $K\coloneqq K(\eta_{\mathcal{T}})$ and $S\coloneqq\{1,\dots,K\}$. Decompose
$Z_{n}=U_{S}+R,\qquad U_{S}\coloneqq\sum_{k\in S}\hat{\lambda}_{k}\,\chi^{2}_{1,k},\qquad R\coloneqq\sum_{k\notin S}\hat{\lambda}_{k}\,\chi^{2}_{1,k}.$
Then $U_{S}$ and $R$ are independent. On $\mathcal{E}_{n}(P)$ we have $\sum_{k>K}\hat{\lambda}_{k}\leq a_{\mathcal{T}}/8$ (as established in the main text from the definition of $\mathcal{E}_{n}(P)$ and the choice $\eta_{\mathcal{T}}\leq a_{\mathcal{T}}/16$), hence
$\operatorname{\mathbb{E}}[R\mid\hat{C}_{n}]=\sum_{k\notin S}\hat{\lambda}_{k}\leq\frac{a_{\mathcal{T}}}{8}.$
With $r_{0}\coloneqq a_{\mathcal{T}}/4$, Markov’s inequality yields
$\mathbb{P}(R\leq r_{0}\mid\hat{C}_{n})\geq 1-\frac{\operatorname{\mathbb{E}}[R\mid\hat{C}_{n}]}{r_{0}}\geq\frac{1}{2}.$
We next obtain a uniform lower bound for the density of $U_{S}$ on a fixed compact interval. For $\hat{\lambda}>0$ define
$h_{\hat{\lambda}}(u)\coloneqq\frac{1}{\hat{\lambda}}\,f_{\chi^{2}_{1}}\left(\frac{u}{\hat{\lambda}}\right)=\frac{1}{\sqrt{2\pi\hat{\lambda}u}}\exp\left(-\frac{u}{2\hat{\lambda}}\right),\qquad u>0,$ (A.17)
and set $h_{\hat{\lambda}}(u)\coloneqq 0$ for $u\leq 0$. Let $I_{x}\coloneqq[a_{\mathcal{T}}/4,\,2b_{\mathcal{T}}]$. Write $U_{S}=X+V$ with
$X\coloneqq\hat{\lambda}_{1}\chi^{2}_{1,1},\qquad V\coloneqq\sum_{k=2}^{K}\hat{\lambda}_{k}\,\chi^{2}_{1,k}.$
Conditionally on $\hat{C}_{n}$, $X$ and $V$ are independent, and $X$ has density $h_{\hat{\lambda}_{1}}$. Hence $U_{S}$ admits a density $g_{S}(\cdot;\hat{\lambda}_{1:K})$ given by the convolution formula
$g_{S}(x;\hat{\lambda}_{1:K})=\operatorname{\mathbb{E}}\left[h_{\hat{\lambda}_{1}}(x-V)\,\middle|\,\hat{C}_{n}\right],\qquad x\in\mathbb{R},$ (A.18)
where $h_{\hat{\lambda}_{1}}(x-V)=0$ whenever $x\leq V$.
On $\mathcal{E}_{n}(P)$, we have $\hat{\lambda}_{1}\in[\lambda_{\star}/2,\,2c_{1}]$ and $\hat{\lambda}_{k}\leq\operatorname{tr}(\hat{C}_{n})\leq 2c_{1}$ for all $k\leq K$. Define the constant
$\underline{h}\coloneqq\inf_{\lambda\in[\lambda_{\star}/2,\,2c_{1}]}\ \inf_{u\in[a_{\mathcal{T}}/8,\,2b_{\mathcal{T}}]}\ h_{\lambda}(u).$
Since $(\lambda,u)\mapsto h_{\lambda}(u)$ is continuous and strictly positive on the compact set $\lambda_{\star}/2,2c_{1}]\times[a_{\mathcal{T}}/8,2b_{\mathcal{T}}]$, we have $\underline{h}>0$.
For $x\in I_{x}$, consider the event $A\coloneqq\{V\leq a_{\mathcal{T}}/8\}$. On $A$ we have $x-V\in[a_{\mathcal{T}}/8,\,2b_{\mathcal{T}}]$, and therefore $h_{\hat{\lambda}_{1}}(x-V)\geq\underline{h}$. Using (A.18),
$g_{S}(x;\hat{\lambda}_{1:K})\geq\underline{h}\,\mathbb{P}(A\mid\hat{C}_{n}).$
To bound $\mathbb{P}(A\mid\hat{C}_{n})$ from below uniformly, note that
$\bigcap_{k=2}^{K}\left\{\hat{\lambda}_{k}\chi^{2}_{1,k}\leq\frac{a_{\mathcal{T}}}{8(K-1)}\right\}\subseteq\left\{V\leq\frac{a_{\mathcal{T}}}{8}\right\}=A.$
By independence of $(\chi^{2}_{1,k})_{k=2}^{K}$ and the bound $\hat{\lambda}_{k}\leq 2c_{1}$,
@@
$\mathbb{P}(A\mid\hat{C}_{n})\geq\prod_{k=2}^{K}\mathbb{P}\left(\chi^{2}_{1}\leq\frac{a_{\mathcal{T}}}{16c_{1}(K-1)}\right)=p_{0}^{K-1},\quad p_{0}\coloneqq\mathbb{P}\left(\chi^{2}_{1}\leq\frac{a_{\mathcal{T}}}{16c_{1}(K-1)}\right)>0.$
Consequently, for all $x\in I_{x}$,
$g_{S}(x;\hat{\lambda}_{1:K})\geq\underline{h}\,p_{0}^{K-1}\eqqcolon m_{S},$
where $m_{S}>0$ depends only on $\lambda_{\star},c_{1},\mathcal{T}$ (through $a_{\mathcal{T}},b_{\mathcal{T}},\eta_{\mathcal{T}}$ and the deterministic index $K(\eta_{\mathcal{T}})$).
Let $f_{n}$ denote the conditional density of $Z_{n}$ given $\hat{C}_{n}$. Since $Z_{n}=U_{S}+R$ with $U_{S}$ independent of $R$ and $U_{S}$ having density $g_{S}(\cdot;\hat{\lambda}_{1:K})$, we have, for all $x\in\mathbb{R}$,
$f_{n}(x)=\int_{[0,\infty)}g_{S}(x-v;\hat{\lambda}_{1:K})\,F_{R}(\mathop{}\!\mathrm{d}v),$
where $F_{R}$ is the conditional distribution function of $R$ given $\hat{C}_{n}$. For any $x\in[a_{\mathcal{T}}/2,\,2b_{\mathcal{T}}]$ and any $v\in[0,r_{0}]$,
$x-v\in\left[a_{\mathcal{T}}/2-r_{0},\,2b_{\mathcal{T}}\right]=\left[a_{\mathcal{T}}/4,\,2b_{\mathcal{T}}\right]=I_{x},$
so $g_{S}(x-v;\hat{\lambda}_{1:K})\geq m_{S}$. Therefore,
$f_{n}(x)\geq\int_{[0,r_{0}]}g_{S}(x-v;\hat{\lambda}_{1:K})\,F_{R}(\mathop{}\!\mathrm{d}v)\geq m_{S}\,\mathbb{P}(R\leq r_{0}\mid\hat{C}_{n})\geq\frac{m_{S}}{2}\eqqcolon m_{0}.$
Finally, for any $\tau\in\mathcal{T}$, the already established bounds $Q_{\tau}(\hat{C}_{n})\in[a_{\mathcal{T}},b_{\mathcal{T}}]$ imply
$Q_{\tau}(\hat{C}_{n})-r_{0}\geq a_{\mathcal{T}}-\frac{a_{\mathcal{T}}}{4}=\frac{a_{\mathcal{T}}}{2},\qquad Q_{\tau}(\hat{C}_{n})+r_{0}\leq b_{\mathcal{T}}+\frac{a_{\mathcal{T}}}{4}\leq 2b_{\mathcal{T}},$
and hence
$\left[Q_{\tau}(\hat{C}_{n})-r_{0},\ Q_{\tau}(\hat{C}_{n})+r_{0}\right]\subseteq\left[\frac{a_{\mathcal{T}}}{2},\ 2b_{\mathcal{T}}\right].$
Thus $f_{n}(x)\geq m_{0}$ holds for all $x$ in the above interval, uniformly over $\tau\in\mathcal{T}$, as claimed. ∎